OpenStudy (anonymous):

what is the simplified form of 5/3y+2 - 4/9y^2-4

OpenStudy (campbell_st):

the denominator in the 2nd fraction \[9y^2 - 4 = (3y -2)(2y+2)\] which needs to be the common denominator the 1st fraction needs to be multiplied by (3y - 2)/(3y -2) so you get \[\frac{5(3y - 2)}{(3y + 2)(3y-2)} - \frac{4}{3y +2)(3y - 2)} = \frac{ 15y - 10 - 4}{(3y +2)(3y -2)}\]

OpenStudy (campbell_st):

just simplify the numerator

OpenStudy (anonymous):

OMG thank you. i was soooooo lost. your incredible. TYTYTY

OpenStudy (campbell_st):

hope it helps